# Rasterized Steered Mixture of Experts for Efficient 2D Image Regression

### v1

@@ V-B Segmentation with modified DBSCAN for Enhanced Denoising
We employ modified DBSCAN [40], a region-based clustering method, to generate these meaningful segments from pixel RGB similarity. More details on this segmentation-based initialization can be found in the cited work.
@@ VI-A Datasets, Metrics, and Baselines
Our experiments focus on two core tasks: image regression and denoising. For both tasks, we use the widely adopted Kodak dataset [41], which comprises 24 high-quality color images at a resolution of 768×512 pixels. This dataset serves as a standard benchmark for evaluating image fidelity and perceptual quality. To further assess the robustness of our method, we additionally include the DIV2K dataset [42], consisting of 100 high-resolution images (1060×768) with diverse natural content. This dataset allows us to examine performance under more varied textures and structures.
@@
We benchmark the R-SMoE model against several state-of-the-art baselines. For image regression, we compare against GS for 2D images (GaussianImage) [5], a recent rasterized RBF method based on Gaussian Splatting (referred to as GS-RBF); the global SMoE (GSMoE) model [21, 19]; and the Radial Basis Function (RBF) approach. While the Radial Basis Function (RBF) framework encompasses any function of the form $\phi(\mathbf{x})=\hat{\phi}(\|\mathbf{x}-\mathbf{u}\|)$, we follow common practice and adopt the Gaussian RBF, defined as $\phi(r)=\exp(-\gamma r^{2})$, consistent with the implementation in [32]. For denoising, we use the well-established BM3D algorithm [43] as the primary baseline.
@@
RBF[32] 2000 27.20 0.7279 0.4156 1746s 0.349s 2.86 4000 876 MB
GSMoE[21] 2000 28.10 0.7662 0.3628 1725s 0.346s 2.89 4000 878 MB
GaussianImage[5] 56 27.59 0.7422 0.3983 98s 2.2ms 449 112 768 MB
@@
RBF[32] 10000 32.38 0.8926 0.2065 7486s 1.51s 0.66 20000 1290 MB
GSMoE[21] 10000 32.98 0.9040 0.1900 7673s 1.53s 0.65 20000 1292 MB
GaussianImage[5] 69 32.82 0.8997 0.1954 104s 2.4ms 415 140 778 MB
@@ VI-B Implementation Details
The R-SMoE model is implemented within the GS-Splat framework [5], extended with specialized CUDA kernels to perform rasterization through a weighted sum of gating functions. The covariance of the 2D Gaussians is parameterized using Cholesky factorization.
@@
Kernel Initialization: Each regression starts with an initial pool of $L$ kernels, randomly initialized with a fixed scaling factor of 5 pixels. The gating mechanism dynamically determines the average number of active kernels per block. For denoising tasks, segmentation boundaries are determined using modified DBSCAN [40], with pixel difference thresholds of 10 and 20.
Training Protocol: All models are optimized using Adam [44] for 10,000 iterations. The learning rates are set to 0.01 for Gaussian centers ($\mu$), 0.001 for covariance matrices ($\Sigma$), and 0.001 for the expert outputs ($m$). The learning rate for $\mu$ decays exponentially to 0.00001. Early stopping is not applied; convergence is empirically determined after 10,000 iterations.
Hardware and Runtime: All experiments are conducted on NVIDIA A4000 GPUs (16 GB) running Ubuntu 20.04 with CUDA 11.8 and PyTorch 2.1. FLOPs are reported following the same per-pixel measurement methodology as [5].
@@
Figure 6: Visualization of image regression results at 6000 kernels for RBF [32], GSMoE [21], GaussianImage [5], and the proposed R-SMoE. Two regions from the original image (left) are cropped and enlarged to highlight differences in edge sharpness and fine-structure reconstruction.
@@
Figure 7: Visual comparison on the DIV2K dataset. R-SMoE achieves comparable perceptual quality to GaussianImage [5] while producing sharper edges and smoother textures with fewer kernels. Error maps are scaled for visibility.
@@ VII-A Image Regression
Table I and Fig. 4 reveal the performance gains of the rasterized SMoE (R-SMoE) model against GaussianImage [5], GSMoE [21, 19], and RBF [32]. Table I summarizes reconstruction quality metrics, while Fig. 4 demonstrates reductions in computational complexity and memory usage. Bold values highlight the best method per group: global methods (above the line) and rasterized methods (below).
@@ VII-B Denoising
Bicubic GaussianImage [5] R-SMoE (Ours) Bicubic GaussianImage [5] R-SMoE (Ours)

### v2

@@ V-B Segmentation with modified DBSCAN for Enhanced Denoising
We employ modified DBSCAN [li_segmentation-based_2023], a region-based clustering method, to generate these meaningful segments from pixel RGB similarity. More details on this segmentation-based initialization can be found in the cited work.
@@ VI-A Datasets, Metrics, and Baselines
Our experiments focus on two core tasks: image regression and denoising. For both tasks, we use the widely adopted Kodak dataset [kodak], which comprises 24 high-quality color images at a resolution of 768×512 pixels. This dataset serves as a standard benchmark for evaluating image fidelity and perceptual quality. To further assess the robustness of our method, we additionally include the DIV2K dataset [DIV2K], consisting of 100 high-resolution images (1060×768) with diverse natural content. This dataset allows us to examine performance under more varied textures and structures.
@@
We benchmark the R-SMoE model against several state-of-the-art baselines. For image regression, we compare against GS for 2D images (GaussianImage) [zhang2024gaussianimage], a recent rasterized RBF method based on Gaussian Splatting (referred to as GS-RBF); the global SMoE (GSMoE) model [verhack_steered_2020, bochinski_regularized_2018]; and the Radial Basis Function (RBF) approach. While the Radial Basis Function (RBF) framework encompasses any function of the form $\phi(\mathbf{x})=\hat{\phi}(\|\mathbf{x}-\mathbf{u}\|)$, we follow common practice and adopt the Gaussian RBF, defined as $\phi(r)=\exp(-\gamma r^{2})$, consistent with the implementation in [ghosh_overview_2001]. For denoising, we use the well-established BM3D algorithm [bm3d_2007] as the primary baseline.
@@
RBF[ghosh_overview_2001] 2000 27.20 0.7279 0.4156 1746s 0.349s 2.86 4000 876 MB
GSMoE[verhack_steered_2020] 2000 28.10 0.7662 0.3628 1725s 0.346s 2.89 4000 878 MB
GaussianImage[zhang2024gaussianimage] 56 27.59 0.7422 0.3983 98s 2.2ms 449 112 768 MB
@@
RBF[ghosh_overview_2001] 10000 32.38 0.8926 0.2065 7486s 1.51s 0.66 20000 1290 MB
GSMoE[verhack_steered_2020] 10000 32.98 0.9040 0.1900 7673s 1.53s 0.65 20000 1292 MB
GaussianImage[zhang2024gaussianimage] 69 32.82 0.8997 0.1954 104s 2.4ms 415 140 778 MB
@@ VI-B Implementation Details
The R-SMoE model is implemented within the GS-Splat framework [zhang2024gaussianimage], extended with specialized CUDA kernels to perform rasterization through a weighted sum of gating functions. The covariance of the 2D Gaussians is parameterized using Cholesky factorization.
@@
Kernel Initialization: Each regression starts with an initial pool of $L$ kernels, randomly initialized with a fixed scaling factor of 5 pixels. The gating mechanism dynamically determines the average number of active kernels per block. For denoising tasks, segmentation boundaries are determined using modified DBSCAN [li_segmentation-based_2023], with pixel difference thresholds of 10 and 20.
Training Protocol: All models are optimized using Adam [diederik2015adam] for 10,000 iterations. The learning rates are set to 0.01 for Gaussian centers ($\mu$), 0.001 for covariance matrices ($\Sigma$), and 0.001 for the expert outputs ($m$). The learning rate for $\mu$ decays exponentially to 0.00001. Early stopping is not applied; convergence is empirically determined after 10,000 iterations.
Hardware and Runtime: All experiments are conducted on NVIDIA A4000 GPUs (16 GB) running Ubuntu 20.04 with CUDA 11.8 and PyTorch 2.1. FLOPs are reported following the same per-pixel measurement methodology as [zhang2024gaussianimage].
@@
Figure 6: Visualization of image regression results at 6000 kernels for RBF [ghosh_overview_2001], GSMoE [verhack_steered_2020], GaussianImage [zhang2024gaussianimage], and the proposed R-SMoE. Two regions from the original image (left) are cropped and enlarged to highlight differences in edge sharpness and fine-structure reconstruction.
@@
Figure 7: Visual comparison on the DIV2K dataset. R-SMoE achieves comparable perceptual quality to GaussianImage [zhang2024gaussianimage] while producing sharper edges and smoother textures with fewer kernels. Error maps are scaled for visibility.
@@ VII-A Image Regression
Table I and Fig. 4 reveal the performance gains of the rasterized SMoE (R-SMoE) model against GaussianImage [zhang2024gaussianimage], GSMoE [verhack_steered_2020, bochinski_regularized_2018], and RBF [ghosh_overview_2001]. Table I summarizes reconstruction quality metrics, while Fig. 4 demonstrates reductions in computational complexity and memory usage. Bold values highlight the best method per group: global methods (above the line) and rasterized methods (below).
@@ VII-B Denoising
Bicubic GaussianImage [zhang2024gaussianimage] R-SMoE (Ours) Bicubic GaussianImage [zhang2024gaussianimage] R-SMoE (Ours)
@@ IX Summary and Conclusion
[Uncaptioned image] Yi-Hsin Li received a Bachelor’s degree in Electrical Engineering from National Chiao Tung University in 2018 and a Master’s degree in Electrical Engineering from National Taiwan University in 2020. In November 2021, she started a double-degree Ph.D. program at the Technical University of Berlin, Germany, with a secondment to Mid Sweden University. She is in her fourth year of Ph.D. research on high-dimensional data compression, focusing on gating networks.
[Uncaptioned image] Mårten Sjöström (M’95, SM’17) received the M.Sc. degree from Linköping University (1992), the Licentiate of Technology degree from the Royal Institute of Technology Stockholm (1998), and the Ph.D. degree from the École Polytechnique Fédérale de Lausanne (2001). He was with ABB (1993-1994) and with CERN (1994-1996), involved with projects on signal processing. In 2001, he joined Mid Sweden University and was appointed Associate Professor (2008) and Full Professor of Signal Processing (2013). He is head of research education in Computer and System Science (2013-) and Computer Engineering (2020-). He is head and founder of the Realistic 3-D Research Group (2007-). He has served as Associate editor for IEEE Transactions on Image Processing (2022-), and for SPIE Journal of Electronic Imaging (2018-2022). He is a board member of High Performance Computing Centre North (2019-). His current research interests include Visual AI, machine learning for multidimensional signal processing and imaging, and system modeling and identification.
[Uncaptioned image] Sebastian Knorr (M’11, SM’19) received the Diploma and Ph.D. degree in Electrical Engineering from the Technical University of Berlin, Germany in 2002 and 2008, respectively. Between 2009 and 2016, he was the CEO/CTO of imcube labs GmbH, Germany. He was Senior Research Scientist and Lecturer at Trinity College Dublin and TU Berlin between 2017 and 2020, respectively. From 2020 to 2024, he was Full Professor at the Ernst-Abbe University of Applied Sciences Jena and is currently Full Professor for Visual Computing at the School of Computing, Communication and Business, HTW Berlin. From 2019 to 2023, he was Associate Editor of the IEEE Transactions of Multimedia, and since 2023, he is Associate Editor of the IEEE Transactions on Image Processing. His research interests include light field imaging, neural rendering, free-viewpoint-video, 3D image processing and 360 video.
[Uncaptioned image] Thomas Sikora (Senior Member, IEEE) is the Director of the Communication Systems Lab at Technische Universität Berlin, Germany. He received the Dipl.-Ing. and Dr.-Ing. degrees in electrical engineering from Bremen University, Bremen, Germany, in 1985 and 1989, respectively. In 1990, he joined Siemens Ltd. and Monash University, Melbourne, Australia, as a Project Leader responsible for video compression research activities in the Australian Universal Broadband Video Codec consortium. Between 1994 and 2001, he was the Director of the Interactive Media Department at the Heinrich Hertz Institute (HHI) Berlin GmbH, Germany. In 2002, he was appointed Full Professor at TU Berlin. Prof. Sikora has been involved in international ITU and ISO standardization activities, as well as in several European research projects, for many years. As Chairman of the ISO-MPEG (Moving Picture Experts Group) video group, he was responsible for the development and standardization of the MPEG-4 and MPEG-7 video algorithms. He is the recipient of an Engineering Emmy Award and the 1996 German ITG Award. He has served as Associate Editor on the editorial boards of several journals, including EURASIP Signal Processing: Image Communication. From 1998 to 2002, he was an Associate Editor of the IEEE Signal Processing Magazine. He is also a past Editor-in-Chief of the IEEE Transactions on Circuits and Systems for Video Technology.